# Learning Mamba as a Continual Learner

### v1

@@ 3.2.2 Regularizing Selectivity of Mamba for Meta-Training
It is non-trivial to meta-learn the continual learner for associating the input and target by seeing a data stream, for both Transformers and attention-free models. The meta-training can be slow to converge or hard to find an optimal solution. We thus consider giving additional guidance in meta-training, by enhancing the association between the query tokens (i.e., testing input) with correlated preceding tokens. During training, for an input ${\mathbf{x}}$ (corresponding to a pair $({\mathbf{x}},y)$) in the stream at the step $12t+$ 1 after $2t$ tokens of $t$ samples, its association relationship with preceding tokens can be represented as ${\mathbf{p}}_{2t+1}=[\mathbbm{1}_{y_{2t+1}}(y_{1}),\mathbbm{1}_{y_{2t+1}}(y_{1}),...,\mathbbm{1}_{y_{2t+1}}(y_{t}),\mathbbm{1}_{y_{2t+1}}(y_{t})]$ with ${\mathbf{p}}\in\{0,1\}^{2t}$, where $\mathbbm{1}_{y}(y^{\prime})$ is an indicator function with $\mathbbm{1}_{y}(y^{\prime})=1,\text{if}~{}y=y^{\prime},~{}\text{and}~{}\mathbbm{1}_{y}(y^{\prime})=0,\text{if}~{}y\neq y^{\prime}$. We hope the meta-learned learner can also identify and use this pattern in CL (i.e., meta-testing).
@@
Transformers maintain the key-value pairs for all samples as the state. For prediction at a step, attention is applied to all the stored keys through a query, retrieving the learned information. As shown in Eq. (1), for the token at step $12t+$ 1, the attention weights/patterns to previous-step tokens can be denoted as ${\mathbf{q}}_{2t+1}^{\text{Trans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}\in{\mathbb{R}}^{2t}$. Note that we omit the normalization terms in attention weights to simplify the presentation. The meta-learning guidance can be applied by encouraging the similarity between ${\mathbf{q}}_{2t+1}^{\text{Trans}}$ and ${\mathbf{p}}_{2t+1}$.
@@
Mamba and other attention-free methods (e.g., Linear Transformer) compress knowledge in a hidden state at each step, as shown in Eq. (3) and (5). Specifically, Mamba applies an input-dependent selection and gating at each step. Although there are no explicit attention weights produced in Mamba, we formulate the regularization for the selectivity of Mamba by bridging the selective SSM (in Eq. (5) and (6)) with linear attention (in Eq. (3)) and the softmax attention (in Eq. (1)). As shown in Eq. (3), Linear Transformer updates the state ${\mathbf{S}}$ (and the normalization term ${\mathbf{G}}$) using kernel-based ${\mathbf{K}}$ and ${\mathbf{V}}$, and performs prediction based on ${\mathbf{Q}}$. Considering that the ${\mathbf{K}}$, ${\mathbf{V}}$, and ${\mathbf{Q}}$ in linear attention share the same meaning as in the softmax attention, we still can obtain ${\mathbf{q}}_{2t+1}^{\text{LNTrans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}$ by storing the ${\mathbf{K}}_{j}$ of intermediate tokens only during training for regularization. By examining the duality relationship between the SSM in Eq. (5) and the formulation of Linear Transformer in Eq. (3(Dao & Gu, 2024), we can identify the connections between the selective parameters, i.e., ${\mathbf{C}}_{t}$ and ${\mathbf{B}}_{t}$, in SSM and query-key embeddings, i.e., ${\mathbf{Q}}_{t}$ and ${\mathbf{K}}_{t}$, in linear attention. Relying on the linear attention as the bridge, we can obtain the associative indicators of Mamba as ${\mathbf{q}}_{2t+1}^{\text{Mamba}}=[{\mathbf{C}}_{2t+1}{\mathbf{B}}_{j}^{\top}]_{j=1}^{2t}$. To regularize the models’ attention or selection behavior in meta-training, for a query sample $({\mathbf{x}},y)$ in a sequence, we apply a selectivity regularization:
$\ell_{\text{slct}}(({\mathbf{x}},y))=\text{KL}({\mathbf{p}}_{\text{idx}(({\mathbf{x}},y))},{\mathbf{q}}^{*}_{\text{idx}(({\mathbf{x}},y))}),$ (8)
where $\text{idx}()$ indicates the step of the token ${\mathbf{x}}$, * indicates the arbitrary model, and KL divergence is used to minimize the difference between model’s association pattern and the ground truth. Note that this regularization and maintained intermediate components are not necessary in inference. We apply this regularization to MambaCL and other sequence prediction models (weighted by a scalar $\lambda$) together with the MCL objective in Eq. (7), which improves the meta-training stability and convergence for all models.
@@ 4 Experiments and Analyses
Experimental setup. To evaluate the performance of various architectures across multiple types of tasks, we conducted a series of experiments. Firstly, we divided one dataset into multiple tasks, typically with each task representing a distinct class within the dataset. We distributed these tasks into two non-overlapping sets, i.e., meta-training and meta-testing. The construction of CL episodes for both meta-groups follows the same procedure: for each CL episode, we randomly select $K$ distinct tasks. $K$ is set as 20 by default. We also investigated scenarios with different of $K$ values. By default, each task in both the training and testing sequences includes five samples (5-shot). Additionally, involving fewer and more shots were also explored to further assess adaptability and learning efficiency.
Datasets. We conduct experiments across various datasets: general image classification tasks included Cifar-100 (Krizhevsky & Hinton, 2009), ImageNet-1K (Russakovsky et al., 2015), ImageNet-R (Russakovsky et al., 2015), MS-Celeb-1M (Celeb) (Guo et al., 2016),CASIA Chinese handwriting (Casia) (Liu et al., 2011), and Omniglot (Lake et al., 2015); fine-grained recognition tasks involved CUB-200 (Wah et al., 2011), Stanford Dogs (Khosla et al., 2011), Stanford Cars (Krause et al., 2013), and FGVC-Aircraft (Aircraft) (Maji et al., 2013); the large domain shift tasks featured (Peng et al., 2019); and regression tasks consisted of sine wave reconstruction (sine), image rotation prediction (rotation), and image completion (completion).
Implementation details. We conduct our main experiments on a single NVIDIA A100 GPU. We repeated each experiment five times and reported the mean and standard deviation of these runs. Results are reported upon convergence on the meta-training set. The batch size is set to 16, and the Adam optimizer is applied. We set the initial learning rate to $1\times 10^{-4}$, with decays of $0.5$ every $10,000$ steps. For all models, we ensure a consistent setup to enable fair comparisons and make sure all models achieve satisfactory results, with additional details provided in Sec. B. Specifically, for experiments involving training from scratch, we adopt the settings from (Lee et al., 2023) to maintain fairness. For the networks built on pre-trained models, we use the OpenAI/CLIP-ViT-B16 (Radford et al., 2021; Ilharco et al., 2021) as our image encoder, with its parameters frozen during training and an additional trainable linear projector.
4.1 Experimental Results and Analyses
In our experiments, we assess several models including OML (Javed & White, 2019), Vanilla Transformer (Vaswani et al., 2017), Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), and our MambaCL. OML serves as a conventional SGD-based meta-continual learning baseline, featuring a two-layer MLP prediction network on top of a meta-learned encoder. Transformers exhibit advanced sequence modeling capabilities, but they may not be optimal for (CL) due to computational inefficiencies and the broad objectives associated with CL. To enhance efficiency, Linear Transformer and Performer utilize kernel methods and linear operations to approximate softmax attention, which maintain a constant hidden state size and exhibit linear computational complexity. All transformer models share a similar structure, each with 4 layers and 512 hidden dimensions. Mamba is an attention-free model optimized for efficiently processing long sequences with near-linear computational demands. Our Mamba Learner also utilizes 4 layers and 512 hidden dimensions, facilitating comparison with the transformer models, yet it features significantly fewer parameters.
Table 1: Classification accuracy (%) across 20-task 5-shot MCL, training from the scratch on general image classification tasks. The best and second best performances are indicated in red and blue, respectively.
@@ 4.1 Experimental Results and Analyses
General image classification tasks. Table 1 and 2 present comparative performance analyses of different architectures on several general image classification tasks, initiating training from scratch and extracting image representation based on a pre-trained model, respectively. In Table 1, within the CIFAR-100 datasets, all methods suffer from substantial meta-overfitting, as evidenced by the large gap between meta-training and meta-testing scores. This may be attributed to the lower task (class) diversity. In Table 2, the results for continual learners built on pre-trained models exhibit similar trends in CIFAR-100 and ImageNet-R. Furthermore, Mamba demonstrates superior performance compared to other methods in these scenarios, underscoring its robustness against overfitting. On larger datasets such as ImageNet-1K, Casia, and Celeb, Mamba performs on par with or surpasses transformers. Without losing generality, we use the pre-trained image representations for our experiments by default.
@@
Fine-grained recognition tasks. Table 3 presents a performance comparison of different architectures on fine-grained recognition datasets. In fine-grained datasets, where only subtle differences exist between classes (e.g., the CUB-200 dataset, which contains 200 bird subcategories), models need capture global information across the entire training episode to distinguish these fine-grained differences. Mamba outperforms other models across these datasets, potentially due to its robustness to capture subtle inter-class distinctions.
@@ 4.2 Generalization Analyses
We hope a meta-learned learner has the ability to be generalized to unseen scenarios. We conduct generalization analyses for Transformer models and Mamba in scenarios involving generalization to longer untrained sequence lengths, larger domain shifts, and sensitivity to the noise inputs, for meta-testing. Additionally, to analyze the behaviors of these models, we visualize the attention weights of Transformers and the associative weights of Mamba to demonstrate their attentions and selectivity patterns in Sec. D.
@@
Figure 3: Generalization Analysis on ImageNet-1K, meta-trained on 20-task 5-shot MCL: (a) meta-testing on varying number of tasks (5-shot); (b) meta-testing on varying number of shots (20-task); (c) meta-testing on varying inputs noise intensity level.
Generalization to different stream length. To effectively address episodes of continual learning of indefinite length, the learning algorithm should demonstrate the capability to generalize beyond the sequence lengths observed during meta-training. We conducted length generalization experiments on ImageNet-1K, training vanilla Transformers, linear Transformers, and Mamba on 20-task 5-shot MCL, each with a vocabulary of 200 tokens. The length of a continual learning episode is calculated as $2\times\text{tasks}\times\text{shots}+1$.
Meta-testing on different numbers of tasks. Fig. 3 shows the performance of the three models meta-trained on a 20-task, 5-shot setup, evaluated during meta-testing across varying numbers of tasks while keeping a constant shot number of 5. Both the Transformer and Linear Transformer suffer significant performance degradation when meta-testing at untrained episode lengths, even for simpler tasks such as the 10-task, 5-shot configuration. Mamba’s meta-testing performance on the 10-task setup is better relative to the meta-trained 20-task setup, and the performance degradation is relatively mild compared to transformers as the number of tasks gradually increases.
Meta-testing on different number of shots. In Fig. 3, we evaluate the performance of three models meta-trained on a 20-task, 5-shot setup, evaluated during meta-testing across varying numbers of shots while maintaining a constant task count of 20. Both the vanilla Transformer and linear Transformer exhibit significant performance degradation, likely due to overfitting the 20-task, 5-shot pattern. However, Mamba experiences only about a 10% performance degradation when the meta-testing shot number reaches 50, which is ten times the meta-training episode length. Fig. 3 and 3 demonstrate Mamba’s robustness in length generalization.
Results and analyses on larger domain shift. We explore a larger domain shift scenario using the DomainNet dataset (containing 6 different domains) to further evaluate model generalization to unseen input distributions, with one domain reserved for meta-testing and the remaining domains for meta-training, which represents a more realistic setting. The experimental results are presented in Table 4. Overall, these models demonstrate the capability to handle large domain shift scenarios. Mamba performs on par with or surpasses Transformer models across various target domains, benefiting from the potentially better generalization ability from smaller-size model with less over-fitting possibility. Vanilla Transformers perform well when the targets are real images or paintings. Mamba excels particularly in the Quickdraw domain, which exhibits larger differences compared to other domains. This performance may be attributed to Mamba’s robustness in processing inputs with larger deviations from the training distribution.
Table 4: Classification accuracy (%) across 20-task 5-shot MCL on DomainNet dataset. (inf,pnt,qdr,rel,skt→clp denotes meta-testing on Clipart domain, and the remaining domains used for meta-training. clp: clipart, inf: infograph, pnt: painting, qdr: quickdraw, rel: real, skt: sketch.)
@@
Method Casia Celeb Sine Rotation Completion
Sensitivity to the noisy inputs. To evaluate the sensitivity of different models to noisy inputs, we conduct experiments on meta-trained 20-task, 5-shot MCL models using the ImageNet-1K dataset. Within each meta-testing episode, we apply noise to the input embeddings ${\mathbf{x}}_{i}$ of five randomly selected samples. We add the noise following Gaussian distributions characterized by a mean ($\mu$) of 00 and a standard deviation ($\sigma$) that ranges from 00 to $10$. As depicted in Fig. 3, the vanilla transformer and linear transformer suffer significant performance degradation. In contrast, Mamba demonstrates robust performance when processing inputs with high levels of noise.
4.3 Traning on Longer Episodes
We conducted experiments to meta-train the models on longer episodes across both classification and regression tasks. Table 5 demonstrates that Mamba continues to perform comparably to Transformer, and significantly outperforms SGD-based approaches (OML).

### v4

@@ 3.2.2 Selectivity Regularization for Meta Training
During training, for an input ${\mathbf{x}}$ (corresponding to a pair $({\mathbf{x}},y)$) in the stream at the step $12t+$ 1 after $2t$ tokens of $t$ samples, its association relationship with preceding input tokens can be represented as ${\mathbf{p}}_{2t+1}=[\mathbbm{1}_{y_{2t+1}}(y_{1}),\mathbbm{1}_{y_{2t+1}}(y_{1}),...,\mathbbm{1}_{y_{2t+1}}(y_{t}),\mathbbm{1}_{y_{2t+1}}(y_{t})]\in\{0,1\}^{2t}$, where $\mathbbm{1}_{y}(y^{\prime})$ is an indicator function with $\mathbbm{1}_{y}(y^{\prime})=1,\text{if}~{}y=y^{\prime},~{}\text{and}~{}\mathbbm{1}_{y}(y^{\prime})=0,\text{if}~{}y\neq y^{\prime}$. We hope the meta-learned learner can also identify and use this pattern in CL (i.e., meta-testing).
@@
Transformers store key-value pairs for all samples as state. At each step, explicit attention retrieves learned information by querying all stored keys. As shown in Eq. (1), for the token at step $12t+$ 1, the attention weights/patterns to previous-step tokens can be denoted as ${\mathbf{q}}_{2t+1}^{\text{Trans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}\in{\mathbb{R}}_{\geq 0}^{2t}$. Note that we omit normalization terms in attention weights to simplify the presentation. With explicit attention, meta-learning guidance can be directly applied for Transformers to encourage similarity between ${\mathbf{q}}_{2t+1}^{\text{Trans}}$ and ${\mathbf{p}}_{2t+1}$.
Mamba and other attention-free methods (e.g., Linear Transformer) compress knowledge in a hidden state at each step, as shown in Eq. (2) and (3). Mamba, in particular, applies input-dependent selection and gating at each step. Although Mamba does not generate explicit attention weights, we formulate a regularization for its selectivity by bridging the selective SSM (Eqs.(3) and (4)) with linear attention and softmax attention in Transformers.
@@
As shown in Eq. (2), Linear Transformer updates the state ${\mathbf{S}}$ (and the normalization term ${\mathbf{G}}$) using kernel-based ${\mathbf{K}}$ and ${\mathbf{V}}$, and performs prediction based on ${\mathbf{Q}}$. Considering that the ${\mathbf{K}}$, ${\mathbf{V}}$, and ${\mathbf{Q}}$ in linear attention share the same meaning as in the softmax attention, we still can obtain ${\mathbf{q}}_{2t+1}^{\text{LNTrans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}$ by storing the ${\mathbf{K}}_{j}$ of intermediate tokens only during training for regularization. By examining the duality relationship between the SSM in Eq. (3) and the formulation of Linear Transformer in Eq. (2[9], we can identify the connections between the selective parameters, i.e., ${\mathbf{C}}_{t}$ and ${\mathbf{B}}_{t}$, in SSM and query-key embeddings, i.e., ${\mathbf{Q}}_{t}$ and ${\mathbf{K}}_{t}$, in linear attention. Relying on the linear attention as the bridge, we can obtain the associative indicators of Mamba as ${\mathbf{q}}_{2t+1}^{\text{Mamba}}=[{\mathbf{C}}_{2t+1}{\mathbf{B}}_{j}^{\top}]_{j=1}^{2t}$. To regularize the models’ selection behavior in meta-training, for a query sample $({\mathbf{x}},y)$ in a sequence, we apply a selectivity regularization: $\ell_{\text{slct}}(({\mathbf{x}},y))=\text{KL}({\mathbf{p}}_{\text{idx}(({\mathbf{x}},y))},{\mathbf{q}}^{*}_{\text{idx}(({\mathbf{x}},y))})$, where $\text{idx}()$ indicates the step of the token ${\mathbf{x}}$, * indicates the arbitrary model, and KL divergence is used to minimize the difference between the model’s association pattern and the ground truth. Note that this regularization and maintained intermediate components are not necessary in inference. We apply this regularization to MambaCL and other sequence prediction models (weighted by a scalar $\lambda$) together with the MCL objective in Eq. (5) to enhance the meta-training stability and convergence. Further details are provided in Sec. 6, Sec. C.1, and Sec. C.3 of the Appendix, where we report the $\lambda$ values used in our selectivity regularization loss and present the training loss curves with and without selectivity regularization for different models.
@@ 4 Experiments and Analyses
We evaluate different models for MCL and analyze their behavior across various scenarios. Our meta-learning episodes primarily involve classification tasks but also include regression (Sec. 4). Unlike previous studies focused on basic settings, we conduct experiments in both standard MCL setups (Sec. 4) and more diverse configurations (Sec. 4.2).
Experimental setup. For obtaining MCL tasks, given a classification dataset, we firstly divided it into multiple tasks, with each task representing a distinct set of classes. We divide these tasks into two non-overlapping sets, i.e., meta-training and meta-testing. CL episodes in both groups are constructed similarly – each episode randomly, where each episode randomly selects $K$ distinct tasks, with $K$ set to 20 by default. We also explore scenarios with varying $K$ values (Sec. 4.2). By default, each task in both training and testing sequences contains five samples (5-shot). We also explore varying shot numbers to further evaluate adaptability and learning efficiency (Sec. 4.2). For implementation details, please refer to the Appendix Sec. B.
Datasets. We conduct experiments on various datasets, including: general image classification tasks included Cifar-100 [30], ImageNet-1K [53], ImageNet-R [53], MS-Celeb-1M (Celeb) [19], CASIA Chinese handwriting (Casia) [36], and Omniglot [31]; fine-grained recognition tasks involved CUB-200 [62], Stanford Dogs [10], Stanford Cars [29], and FGVC-Aircraft (Aircraft) [40]; the large domain shift tasks featured [46]; and regression tasks consisted of sine wave reconstruction (sine), image rotation prediction (rotation), and image completion (completion). Additional details are provided in the Appendix Sec. A.
Table 1: Classification accuracy (%) across 20-task 5-shot MCL, training from scratch on general image classification tasks. Best and second best rsults are highlighted in red and blue, respectively.
@@
4.1 Experimental Results and Analyses
We evaluate several models, including OML [25], Vanilla Transformer [65], Linear Transformer [27], Performer [8], and our proposed MambaCL. OML serves as an SGD-based meta-continual learning baseline with a two-layer MLP atop a meta-learned encoder. While Transformers excel at sequence modeling, their computational inefficiency and broad CL objectives limit their suitability for continual learning. Linear Transformer and Performer address this by approximating softmax attention via kernel methods, achieving linear complexity with fixed hidden state size. All transformer variants use 4 layers with 512 hidden dimensions. Mamba, an attention-free model optimized for long sequences, also adopts this architecture in our MambaCL, but with significantly fewer parameters.
General image classification tasks. Table 1 and 2 provide comparative analyses of the performance of various architectures across a range of general image classification tasks, initiating training from scratch and utilizing image representations derived from a pre-trained, respectively. In Table 1, within the CIFAR-100 datasets, all methods suffer from substantial meta-overfitting, as evidenced by the large gap between meta-training and meta-testing scores. This may be attributed to the lower task diversity. In Table 2, continual learners based on pre-trained models show consistent trends on CIFAR-100 and ImageNet-R, with Mamba outperforming other methods, highlighting its robustness against overfitting. On larger datasets like ImageNet-1K, Casia, and Celeb, Mamba matches or exceeds transformer performance. We default to using pre-trained image representations for our experiments to maintain generality. To better understand its behavior, we visualize the associative weights of our meta-trained Mamba Learner in Fig. 3. For further details on the visualization analysis of Transformers and Mamba, please refer to the Appendix Sec. D.
@@ 4.1 Experimental Results and Analyses
Fine-grained recognition tasks. Table 3 presents a performance comparison of different architectures on fine-grained recognition datasets. In fine-grained datasets characterized by subtle inter-class differences (e.g., the CUB-200 dataset, which contains 200 bird subcategories), models need to integrate global information throughout the training episode to identify these nuances effectively. Mamba outperforms other models across these datasets, potentially attributable to its robustness to capture subtle inter-class distinctions.
Method Casia Celeb Sine Rotation Completion
Training on Longer Episodes. We conducted experiments to meta-train the models on longer episodes across both classification and regression tasks. Table 4 demonstrates that Mamba continues to perform comparably to Transformer, and significantly outperforms SGD-based approaches (OML).
Figure 3: Final-layer associations in a 20-task 5-shot meta-testing episode of MambaCL (meta-trained on 20-task 5-shot MCL). All three visualizations share the same training episode (shots $0^{th}\!\!-\!99^{th}$), with test queries at the $100^{th}$ shot from the $1^{st}$, $9^{th}$, and $18^{th}$ tasks, aligning with the $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ training shots. The red box highlights each query and its corresponding examples. See Appendix Sec. D for additional visualizations comparing Transformer attention and Mamba’s associative selectivity.
@@ 4.2 Generalization Analyses
We hope a meta-learned learner has the ability to be generalized to unseen scenarios. To assess this, we conduct generalization analyses for Transformer models and Mamba in conditions involving longer untrained sequence lengths, significant domain shifts, and sensitivity to noisy inputs during meta-testing. Additionally, to understand the behavior of these models, we visualize the attention weights of Transformers and the associative weights of Mamba, illustrating their attention mechanisms and selectivity patterns in the Appendix Sec. D.
Generalization to varying sequence length. To address continual learning episodes of indefinite length effectively, the learning algorithm should demonstrate the capability to generalize beyond the sequence lengths encountered during meta-training. We conducted length generalization experiments on ImageNet-1K, training vanilla and linear Transformers, and Mamba on 20-task 5-shot MCL, each with a vocabulary of 200 tokens. The length of a continual learning episode is calculated as $2\times\text{tasks}\times\text{shots}+1$.
(a) Meta-testing on varying numbers of tasks. Fig. 5 illustrates the performance of the three models, meta-trained on a 20-task 5-shot setup and evaluated during meta-testing across varying numbers of tasks while keeping a constant shot number of 5. Both the vanilla Transformer and Linear Transformer suffer significant performance degradation during meta-testing at untrained episode lengths, even in configurations such as the 10-task 5-shot setup. Mamba demonstrates superior performance in the 10-task setup compared to the meta-trained 20-task setup, and its performance degradation is relatively mild compared to transformers as the number of tasks increases.
(b) Meta-testing on varying numbers of shots. In Fig. 5, we assess the performance of three models meta-trained in a 20-task 5-shot setup and evaluated during meta-testing across varying shots numbers while keeping a constant task number of 20. Both the vanilla Transformer and linear Transformer suffer significant performance degradation, likely due to overfitting the 20-task 5-shot pattern. In contrast, Mamba exhibits only about a 10% performance degradation when the number of shots in meta-testing increases to 50, which is ten times the length of the meta-training episodes. Fig. 5 and 5 highlight Mamba’s robustness in adapting to longer sequence lengths (length generalization).
@@
Figure 4: Generalization analysis on ImageNet-1K, with meta-training on 20 tasks (5-shot), meta-testing on: (a) varying tasks (5-shot), (b) varying shots (20-task), and (c) varying input noise levels (20-task 5-shot).
(a) $\lambda$
(b) SSM state size
Figure 5: Ablation studies: (a) varying $\lambda$ in training loss; (b) varying SSM state size.
Table 5: Classification accuracy (%) across 20-task 5-shot MCL on DomainNet dataset. (inf,pnt,qdr,rel,skt→clp denotes meta-testing on the Clipart domain, and the remaining domains used for meta-training. clp: clipart, inf: infograph, pnt: painting, qdr: quickdraw, rel: real, skt: sketch.)
@@
Results and analyses on larger domain shift. We investigate a larger domain shift scenario using the DomainNet dataset, which includes six distinct domains, to further assess model generalization to unseen input distributions, a setting that mirrors real-world conditions. One domain is held out for meta-testing, while the others are used for meta-training. The experimental results are detailed in Table 5. Overall, these models demonstrate the capability to handle large domain shift scenarios effectively. Mamba performs on par with or surpasses transformers across diverse target domains, benefiting from potentially enhanced generalization capabilities associated with its smaller size and reduced susceptibility to overfitting. Vanilla Transformers perform well with the targets are real images or paintings. Mamba excels in the quickdraw domain, which presents greater variances from other domains. This performance advantage may be ascribed to Mamba’s robustness in handling inputs that deviate significantly from the training distribution.
Sensitivity to the noisy inputs. To assess the sensitivity of different models to noisy inputs, we conduct experiments on meta-trained 20-task 5-shot MCL models using the ImageNet-1K dataset. During each meta-testing episode, we apply noise to the input embeddings ${\mathbf{x}}_{i}$ of five randomly selected samples, adhering to Gaussian distributions with a mean ($\mu$) of 00 and a standard deviation ($\sigma$) varying from 00 to $10$. As illustrated in Fig. 5, the vanilla transformer and linear transformer suffer significant performance degradation. In contrast, Mamba demonstrates robust performance, effectively handling inputs with increased noise levels/strength.